A motorboat travels 500mi in 5 hours going upstream and 1170mi in 9 hours going downstream. What is the rate of the boat in still water an what is the rate of the current?
Answer by lwsshak3(11628)

A motorboat travels 500mi in 5 hours going upstream and 1170mi in 9 hours going downstream. What is the rate of the boat in still water an what is the rate of the current?
**
let x = speed of boat in still water
let c = speed of current
x+c =speed of boat downstream
x-c =speed of boat upstream
travel time=distance/speed
..
500/(x-c)=5
1170/(x+c)=9
..
5x-5c=500
9x+9c=1170
..
45x-45c=4500
45x+45c=5850
90x=10350
x=115 mph
..
-45c=-1350
c=15 mph
..
ans:
rate of speed of boat in still water=115 mph
rate of speed of current=15 mph